OpenStudy (anonymous):

find the percentages 3/4% of 975

OpenStudy (anonymous):

do 0.0075x975

OpenStudy (anonymous):

kay

OpenStudy (anonymous):

A percentage is a part out of 100, so 3/4% = 3/400. -> (3/400)*975 = 117/16